You have 450,000 INR saved up for your vacation. In preparation for the trip you want to exchange your Indian Rupee INR for EUR so you get the following quotes:

Spot rate on the EUR/USD cross rate

1.32

Spot rate on the INR/USD cross rate

65.0

a. What is the INR/EUR cross rate?

b. How many EUR will you obtain for your INR?
Answer #1

a) INR/EUR cross rate = ( INR / USD) * (USD / EUR)

                              = 65 * (1/1.32)

                              = INR 49.24242 per EUR

B)EUR obtained = 450000/49.24242

                  = 9138.4623